OpenStudy (anonymous):

if cos(pi/4) is sqrt2/2 i thought sec(pi/4) is 2/sqrt 2,but its just sqrt 2 . can some one please explain.

OpenStudy (bahrom7893):

Okay: \[2/\sqrt{2} = (\sqrt{2})^2/\sqrt{2} = \sqrt{?}\]

OpenStudy (bahrom7893):

sqrt(2), i meant lol not question mark

OpenStudy (anonymous):

2/sqrt(2) and sqrt(2) are the same. Just rationalize as bahrom did.

OpenStudy (anonymous):

oh i see so i have to rationalize it. thank you guys